​If output is increased from 47 units to 48 units per day


If output is increased from 47 units to 48 units per day, and total costs increase from $1,231 to $1,257 per day, the marginal cost per unit is 


a) $1,257 

b) $26.00 

c) $26.19 

d) $26.25 
Answer #1

MC = TC/Q  

TC = 1257 - 1231 = 26

Q = 48 - 49 = 1

MC =  TC/Q  

MC = 26/1

MC = 26

So marginal cost per unit is $ 26

Thus, (b) is correct option